HETAV PATEL

Grade 11,

10 different books and 2 different pen are given to 3 boys so that each get equal no. of things.Probability that same boy doesn't receive both pens is__________ ans:8/11

10 different books and 2 different pen are given to 3 boys so that each get equal no. of things.Probability that same boy doesn't receive both pens is__________

ans:8/11

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear shefali

probability of same boy doesn't receive both pens = 1 - probabilty of any  boy  receive both pens

since all the student recieves same number of article.so each one get (10+2)/3 =4         

number of ways in which 10 books and 2 pen is distributed to 3 student is =12C4 8C4 4C4

 now find the number of ways such that any boy recieve the both pen...

select one boy from 3 and give both pen to him in = 3C1 * 1 ways

 rest two article can be given to this boy in = 10C2 ways 

  other 2 boys will get article in 8C4 *4C4 ways

so number of ways such that any boy recieve the both pen= 3C1 * 1*10C2 *8C4 *4C4

so required probability = 1- 3C1 * 1*10C2 *8C4 *4C4 / 12C4 8C4 4C4

                                          =8/11

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin